OpenStudy (anonymous):

Aki's bicycle Designs has determined that when x hundred bicycles are built, the average cost per bicycle is given by C(x)=0.1x^2-1.3x+5.205, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?

OpenStudy (anonymous):

minimize by finding the vertex using \[-\frac{b}{2a}=-\frac{-1.2}{2\times .1}=6.5\]

OpenStudy (anonymous):

so answer is 650 because these are in hundreds

OpenStudy (anonymous):

C(6.5)=0.1(6.5)^2-1.3(6.5)+5.205 is needed to find the other coordinate